# Rethinking Visual Layer Selection in Multimodal LLMs

### v1

@@ 9.3 Evaluation Metrics
To facilitate evaluation, we use lmms-eval as our primary evaluation tool. For the MMVet dataset, evaluations must be conducted on the official platform by uploading the necessary data. Regarding the CVBench 3D tasks, where models generally exhibit weaker instruction-following performance, we employ the DeepSeek API as the judge. This tool provides results consistent with GPT-4o but is significantly more cost-effective.
9.4 The impact of LLMs
@@
9.5 The impact of Data Scale
The conclusion remains valid across different data scales. Under the 737k data scale, half of the subtasks in the MME dataset achieve optimal performance using the penultimate layer. However, for tasks like Count, Position, and Existence, the middle visual representation layer (Layer 18) demonstrates either superior or comparable performance. Similarly, results under the 1M data scale also show that half of the optimal performances are achieved on layers other than the penultimate one. The results for SEEDBench subtasks, as presented in Table 9, further support this observation. At the 737k data scale, Layer 18 from the middle representation space achieves the best performance on 5 out of 9 subtasks, while the penultimate layer excels in only 3 subtasks. Likewise, under the 1M data scale, half of the subtasks continue to achieve their best performance on layers other than the penultimate one. These findings consistently demonstrate across varying data scales that shallow and middle layers have the potential to outperform deep layers in certain scenarios.
@@ 9.5 The impact of Data Scale
10 Guidelines on Alignment

### v2

@@ B.2 Evaluation Metrics
To facilitate evaluation, we use lmms-eval as our primary evaluation tool. Evaluations must be conducted on the official platform for the MMVet dataset by uploading the necessary data. Regarding the CVBench 3D tasks, where models generally exhibit weaker instruction-following performance, we employ the DeepSeek API as the judge. This tool provides results consistent with GPT-4o but is significantly more cost-effective.
B.3 The impact of LLMs
@@
B.4 The impact of Data Scale
The conclusion remains valid across different data scales. Under the 737k data scale, half of the subtasks in the MME dataset achieve optimal performance using the penultimate layer. However, for tasks like Count, Position, and Existence, the middle visual representation layer (Layer 18) demonstrates either superior or comparable performance. Similarly, results under the 1M data scale also show that half of the optimal performances are achieved on layers other than the penultimate one.
The results for SEEDBench subtasks, as presented in Table 10, further support this observation. At the 737k data scale, Layer 18 from the middle representation space achieves the best performance on 5 out of 9 subtasks, while the penultimate layer excels in only 3 subtasks. Likewise, under the 1M data scale, half of the subtasks continue to achieve their best performance on layers other than the penultimate one. These findings consistently demonstrate across varying data scales that shallow and middle layers have the potential to outperform deep layers in certain scenarios.
B.5 Layer Selections and Feature Fusion
| Models | General | OCR | Vision-Centric | Hallu | |||||||||
In the shallow layer, we select layers 1 and 3 as representatives. Layer 1, being the most chaotic, primarily captures early-stage visual features, while layer 3 is still in a chaotic state but performs relatively well. In the middle layer, we choose layers 18 and 17, as they achieve the first and second-best performance within this representation space. For the deep layer, we select layer 23, as it demonstrates the highest overall performance. The baseline configuration considers only layer 23 as the visual representation.
Full version of comparison study as shown in Figure 8. We evaluate four state-of-the-art fusion methods, including three from DenseConnector (DC) Yao et al. (2024) and one from MMFuser Cao et al. (2024). As shown in Table 3, our method outperforms STI, SCI, and DCI on 8, 7, and 7 benchmarks, respectively. Compared to MMFuser Cao et al. (2024), our approach demonstrates superior performance on 8 benchmarks. These results highlight the significant potential of visual feature fusion strategies in enhancing MLLMs and offering guidance for developing future fusion strategies.
@@ B.5 Layer Selections and Feature Fusion
Appendix C Detailed Discussion
C.1 Alignment Guideline
@@
C.2 Comparison with Prior Work
Although Liu et al. (2023b); Yao et al. (2024); Cao et al. (2024); Li et al. (2026) have briefly considered features from different layers, none have conducted a systematic or in-depth analysis of layer selection or the functional roles of visual layers in MLLMs.
| Model | MMB | SEEDB | GQA | OCRB | RefCOCO |
| Baseline (Vicuna-7B) | 65.1 | 65.2 | 63.2 | 320 | 52.3 |
| DenseConnector Yao et al. (2024) | 63.3 | 66.6 | 62.9 | 323 | - |
| MMFuser Cao et al. (2024) | 67.5 | 60.8 | 62.8 | - | - |
| Instruction-Guided Fusion Li et al. (2026) | 66.9 | 68.3 | 63.1 | - | - |
| Ours | 67.4 | 67.7 | 63.5 | 325 | 54.0 |
Liu et al. (2023b) presents a limited comparison between the penultimate and final layers on ScienceQA, without broader analysis. In contrast, we generalize this observation across over 10 benchmarks and various model and data scales, and offer an interpretation in Section 4.1 grounded in CLIP’s contrastive learning objective.
Yao et al. (2024) uses attention maps to suggest that different layers attend to different regions of interest. However, such analysis is inherently qualitative and subjective, lacking a principled explanation of why these differences arise or how they relate to downstream tasks.
Cao et al. (2024) proposes a shallow-to-deep fusion strategy but does not explain the underlying rationale for selecting specific layers, and its choices are based on empirical heuristics (e.g., 1/3, 2/3, 3/3), with no analysis of representations itself.
While Li et al. (2026) focuses on fusion design, it lacks consideration of the role of visual layers in MLLMs. Our work provides novel insights, such as the unexpected correlation between inter-layer similarity and downstream performance. Moreover, they analyze only 1/5 of ViT layers, without considering data scale or model size.
In contrast, our work is the first to systematically group ViT layers based on representation similarity, and to establish clear correlations between layer types and multimodal task performance across 60+ tasks and multiple model sizes (1.4B–7B) and data scales (665k-1M).
Appendix D Ethics Statement
This work made limited use of ChatGPT for language polishing and code debugging support. All scientific contributions, analyses, and conclusions are those of the authors.
| Index | MME | MMVet | MMBench | SEEDBench |
| 1 | Code Reasoning | Rec_OCR_Spat_Math | Attribute Comparison | Instance Attribute |
| 2 | Count | Rec_Spat | Attribute Recognition | Instance Location |
| 3 | Numerical Calculation | Rec_OCR_Know_Gen | Physical Property | Instance Interaction |
| 4 | Position | OCR_Spat | Spatial Relation | Text Recognition |
| 5 | Existence | OCR_Know_Spat | Image-Text Understanding | Instance Identity |
| 6 | Scene | Rec | Fine-grained Perception (Cross-instance) | Scene Understanding |
| 7 | OCR | OCR_Gen_Spat | Image Style | Instance Counting |
| 8 | Poster | Rec_Know_Gen | Physical Relation | Spatial Relation |
| 9 | Color | Rec_Know | Image Scene | Visual Reasoning |
| 10 | Celebrity | Rec_OCR_Gen_Spat | Celebrity Recognition | — |
| 11 | Landmark | OCR | Identity Reasoning | — |
| 12 | Artwork | Rec_OCR_Gen | Image Emotion | — |
| 13 | Commonsense Reasoning | Rec_OCR | Image Topic | — |
| 14 | — | — | Natural Relation | — |
| 15 | — | — | Object Localization | — |
| 16 | — | — | OCR | — |
| 17 | — | — | Social Relation | — |
| 18 | — | — | Attribute Recognition | — |
| 19 | — | — | Coarse Perception | — |
| 20 | — | — | Fine-grained Perception (Single-instance) | — |
| 21 | — | — | Relation Reasoning | — |
| 22 | — | — | Action Recognition | — |
| 23 | — | — | Function Reasoning | — |
| 24 | — | — | Future Prediction | — |
| 25 | — | — | Logical Reasoning | — |